A solution contains 20.0g of NaCl in 750ml distilled water is found to have an absorbance of 2:52 at 520nm when the path length is 1cm. Calculate the molar absorptivity with units.

1
Expert's answer
2021-11-15T10:55:15-0500

Molar mass of NaCl= 58.44

20/58.44= 0.342

750/1000= 0.75

Molarity= 0.342/0.75= 0.456mol/L

0.456×520= 237.28